OpenStudy (inowalst):

If the average (arithmetic mean) of a, b, c, and d is equal to the average of a, b, and c, what is d in terms of a, b, and c?

OpenStudy (inowalst):

I got: \[\frac{ a+b+c }{ 3 }\]Since for the mean, its to add up all the numbers and divide the sum by the number of numbers.

OpenStudy (inowalst):

@TheSmartOne

TheSmartOne (thesmartone):

Do if we take what your question gave us we get... \(\sf\Huge \frac{a+b+c+d}{4} = \frac{a+b+c}{3}\)

TheSmartOne (thesmartone):

So we cross multiply and we just solve for d @inowalst

OpenStudy (inowalst):

Oh okay.

OpenStudy (inowalst):

a + b + c. Cancel out?

TheSmartOne (thesmartone):

we can't really do that. Unless it was \(\sf a \times b \times c\)

OpenStudy (inowalst):

Ohh. :/ Okay.

TheSmartOne (thesmartone):

\(\sf\huge 3(a+b+c+d) = 4(a+b+c)\)

OpenStudy (inowalst):

I have a question..

TheSmartOne (thesmartone):

Or we could do it like \(\sf\huge 12 \times(\frac{a+b+c+d}{4}) = 12 \times (\frac{a+b+c}{3})\)

TheSmartOne (thesmartone):

Which ofc gets us back to \(\sf\huge 3(a+b+c+d) = 4(a+b+c)\)

OpenStudy (inowalst):

Ohh okay. I think I found the answer because I have options and you said one of them.

TheSmartOne (thesmartone):

Coolio :)

OpenStudy (inowalst):

Thank you!

OpenStudy (theraggedydoctor):

dot.